# Cosmological Zoom-In Simulation of Odd Radio Circles as Merger-Driven Shocks in Galaxy Groups

### v1

@@ Acknowledgements.
AI, KD and IK acknowledge support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679. LMB is supported by NASA through grant 80NSSC24K0173. The calculations for the hydrodynamical simulations were carried out at the Leibniz Supercomputer Center (LRZ) under the project pn68va. The following software was used for this work: julia (Bezanson et al. 2017), Matplotlib (Hunter 2007), Smac (Dolag et al. 2005a).References
@@
- Gupta et al. (2025) Gupta, N., Norris, R. P., Hayder, Z., et al. 2025, arXiv e-prints, arXiv:2506.08439
@@
- Koribalski et al. (2025) Koribalski, B. S., Dolag, K., Khabibullin, I., et al. 2025, arXiv e-prints, arXiv:2507.11781
@@
- Norris et al. (2021) Norris, R. P., Intema, H. T., Kapińska, A. D., et al. 2021, PASA, 38, e003
@@
Appendix A Rotation measure
Fig. 8 shows the correction to the polarization angle due to Faraday rotation along the line of sight (c.f. Eq. 5) of the simulated halo. We assume a frequency of which corresponds to a wavelength of . The contribution in the central region is clearly dominating, while the outer region at the ORC position is adding insignificant rotation. This is mostly due to the much smaller gas density at these radii. Therefore Faraday rotation of synchrotron emission plays only a minor role in determining the polarization fraction at the shock front (c.f. Sect. 3.5).
Appendix B Entropy features
@@
where and are the local temperature and density of the gas respectively. No coherent global structures are visible. Instead, the system displays a random entropy excess pattern, which is indicative of the turbulent motion in action inside the circumgalactic medium. It is mostly driven by the small satellite structures orbiting within the group halo.

### v2

@@ Acknowledgements.
We thank the anonymous referee for their helpful comments and suggestions, which greatly contributed to the quality of this work. AI, KD and IK acknowledge support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679. LMB is supported by NASA through grant 80NSSC24K0173. The calculations for the hydrodynamical simulations were carried out at the Leibniz Supercomputer Center (LRZ) under the project pn69va and with support of the Computational Center for Particle and Astrophysics (C2PAP). The following software was used for this work: julia (Bezanson et al. 2017), Matplotlib (Hunter 2007), Smac (Dolag et al. 2005a).References
@@
- Gupta et al. (2025) Gupta, N., Norris, R. P., Hayder, Z., et al. 2025, PASA, 42, e097
@@
- Koribalski et al. (2025) Koribalski, B. S., Dolag, K., Khabibullin, I., et al. 2025, submitted to PASA, arXiv:2507.11781
@@
- Norris et al. (2021a) Norris, R. P., Crawford, E., & Macgregor, P. 2021a, Galaxies, 9, 83
- Norris et al. (2021b) Norris, R. P., Intema, H. T., Kapińska, A. D., et al. 2021b, PASA, 38, e003
@@
Appendix A Mach number distribution
We show in Fig. 8 the Mach number distribution for the gas particles which make up the shock structure identified as an ORC in the simulation. For the same particles, we showed the time-dependent radio power and synchrotron emission spectrum in Fig. 4. The three panels represent the distribution with three different weightings: unweighted (top), density weighted (middle) and synchrotron emissivity weighted (bottom). In each case, the distribution is dominated by contributions below , with a tail towards high Mach numbers. The broadness of the bulk distribution varies with the weighting method, while the maximum of the distribution stays relatively constant around .
Appendix B Rotation measure
Fig. 9 shows the correction to the polarisation angle due to Faraday rotation along the line of sight (see Eq. 5) of the simulated halo. We assume a frequency of which corresponds to a wavelength of . While the central region displays significant Faraday rotation, the outskirts at radial distances of about – where also the ORC is situated – have rotation angles close to 0. Depolarisation of the signal from the radio ring is hence only caused by turbulence of the emission centres along the line of sight and not by successive Faraday rotation. This is mostly due to the much smaller gas density at these radii. Therefore, Faraday rotation of synchrotron emission plays only a minor role in determining the polarisation fraction at the shock front (see Sect. 3.5).
Appendix C Entropy features
@@
where and are the local temperature and density of the gas, respectively. No coherent global structures are visible. Instead, the system displays a random entropy excess pattern, which is indicative of the turbulent motion in action inside the circumgalactic medium. It is mostly driven by the small satellite structures orbiting within the group halo.
Appendix D Magnetic field strength
In Fig. 11 we show the distribution of magnetic field strengths inside the simulated ORC. The three magnetic field models employed in this figure were introduced in Sect. 4.1. It is clearly visible how the tail of low field strengths disappears when moving from the simulated magnetic field to either equipartition model, or . The equipartition models also shift the maximum of the distribution towards a higher value, which is then of the order of a G. The turbulent field model – which was employed in the calculation of the synchrotron emissivity throughout the analysis of this paper – produces overall the highest field strengths.
Appendix E Physalis flux map